# Deep Learning and genetic algorithms for cosmological Bayesian inference speed-up

### v1

@@ Abstract
In this paper, we present a novel approach to accelerate the Bayesian inference process, focusing specifically on the nested sampling algorithms. Bayesian inference plays a crucial role in cosmological parameter estimation, providing a robust framework for extracting theoretical insights from observational data. However, its computational demands can be substantial, primarily due to the need for numerous likelihood function evaluations. Our proposed method utilizes the power of deep learning, employing feedforward neural networks to approximate the likelihood function dynamically during the Bayesian inference process. Unlike traditional approaches, our method trains neural networks on-the-fly using the current set of live points as training data, without the need for pre-training. This flexibility enables adaptation to various theoretical models and datasets. We perform simple hyperparameter optimization using genetic algorithms to suggest initial neural network architectures for learning each likelihood function. Once the sufficient accuracy is achieved, the neural network replaces the original likelihood function. The implementation integrates with nested sampling algorithms and has been thoroughly evaluated using both simple cosmological dark energy models and diverse observational datasets. Additionally, we explore the potential of genetic algorithms for generating initial live points within nested sampling inference, opening up new avenues for enhancing the efficiency and effectiveness of Bayesian inference methods.
@@ I Introduction
Bayesian inference is a powerful tool in several scientific fields where it is essential to constrain mathematical models using experimental data. It allows parameter estimation and model comparison. In particular, it is the data analysis technique par excellence in observational cosmology, as it provides a very robust method to obtain valuable statistical information from a theoretical model given a set of observational data. However, a significant disadvantage of Bayesian inference lies in its high computational cost; it requires a considerable number of likelihood function evaluations to generate sufficient samples from the posterior distribution. For example, a small Bayesian inference task could involve thousands of samples and require thousands, or even millions, of likelihood evaluations.
@@
The use of artificial neural networks (ANNs) to approximate the likelihood function can greatly improve the efficiency of Bayesian inference auld2007fast ; graff2012 ; graff2014 ; hortua2020accelerating ; hortua2020parameter ; spurio2022cosmopower ; nygaard2023connect . However, is necessary to have a careful consideration of the trade-off between accuracy and speed, along with quality monitoring of the resulting posterior samples. In addition, neural networks present several drawbacks that must be taken into account to effectively aid in the performance of Bayesian inference:
@@
Previous works using neural networks in cosmological parameter estimation save an amazing amount of computational time training neural networks before the Bayesian inference process spurio2022cosmopower ; hortua2020accelerating ; chantada2023nn ; however, the pre-training time in these cases is expensive and the trained neural networks are only useful for a specific configuration of backgrounds, models, and data sets. For this reason, our work is inspired by BAMBI graff2012 ; graff2014 , and pyBambi pybambi , where their neural networks are trained in real-time to learn the likelihood function, which is subsequently replaced within a nested sampling process. The strength of this approach lies in its ability to train the neural network in real-time and accelerate the Bayesian inference process without restricting a cosmological or theoretical model and specific datasets. In our method, we explore features beyond those of our predecessors, such as parallelism, PyTorch implementation, and hyperparameter tuning. In addition, we exclusively used live points for training to reduce the dispersion of the training dataset and to obtain results with higher accuracy. A criterion was also chosen to initiate our method that serves as a regulator of the trade-off between accuracy and speed. We also implemented an on-the-fly performance evaluation to accept or reject the neural network predictions. In addition, we have conducted a preliminary investigation on the use of genetic algorithms to generate the initial sample of live points on the nested sampling process.
@@ II.1 Bayesian inference
and if the error $\epsilon\sim N(0,C)$ has a normal distribution centered in zero and a covariance matrix $C$, then:
$P(D|\theta)=\frac{1}{(2\pi)^{N/2}|C|^{1/2}}e^{-0.5(D-f(x;\theta))^{T}C^{-1}(D-f(x;\theta))}\;\;,$ (4)
@@ II.2 Nested sampling
NS starts with a specific number $n_{\rm live}$ of random points, termed live points, distributed within the prior volume defined by the constrained prior. These samples are ordered based on their likelihood values. During each iteration, the worst point (with the lowest likelihood value $\mathcal{L}_{\rm worst}$) is removed. A new sample is then generated within a contour bounded by $\mathcal{L}_{\rm worst}$ and with a likelihood, $\mathcal{L}(\theta)>\mathcal{L}_{\rm worst}$. Equation 7 can be simplified as a Riemann sum:
@@
where $\omega_{i}$ is the difference between the prior volume of two consecutive points: $\omega_{i}=X_{i-1}-X_{i}$. Throughout the process, NS retains the population of $n_{live}$ live points, and ultimately consolidates the final set of live points within a region of high probability. Depending on the sampling approach employed from the constrained prior, various nested sampling algorithms exist. For instance, MultiNest feroz2009multinest utilizes rejection sampling within ellipsoids, whereas Polychord handley2015 generates points using slice sampling.
@@ III Machine Learning background
Machine learning is the field of Artificial Intelligence concerning to the mathematical modeling of datasets. Its methods identify inherent properties of datasets by minimizing a target function until it reaches a satisfactory value. IOver the past few years, Artificial Neural Networks (ANNs) have emerged as the most successful type of machine learning models, giving rise to the field of deep learning. On the other hand, genetic algorithms are a special class of evolutionary algorithms, called metaheuristics, facilitating function optimization without derivatives.
@@ III.1 Artificial neural networks
The intrinsic parameters of a neural network are known as hyperparameters, are set before training, and include parameters such as the number of layers and neurons, epochs, and activation functions. Parameters of gradient descent and backpropagation algorithms rumelhart1986learning , like batch size and learning rate, may also be hyperparameters. While some hyperparameters are predetermined, others are adjusted through tuning strategies.
ANNs are valued for their capacity to model large and complex datasets. The Universal Approximation Theorem asserts that an ANN with a single hidden layer and non-linear activation functions can model any nonlinear function hornik1990universal , enhancing its utility for datasets with complex relationships. Though an exhaustive review of ANNs is beyond the scope of this paper, great references exist in the literature nielsen2015neural ; goodfellow2016deep . For a basic introduction to their algorithms and cosmological context, we recommend reading de2022observational .
Below we describe some deep learning techniques utilized in our training:
  • Data scaling. A preprocessing technique involving the transformation of the scale or range of the dataset. This is particularly beneficial for improving the performance of some machine learning algorithms that are sensitive to the scale of the input data.
  • Early stopping. It is a regularization technique that monitors the performance of a model on a validation set during training and stops the training process when the performance on the validation set starts to degrade, indicating overfitting. It helps to prevent overfitting and choose the best weight configuration along the epochs of the training. By stopping the training process early, the generalization performance of the model can be improved, particularly when the training data is limited or noisy.
  • Dynamic learning rate. The purpose of the dynamic learning rate is to improve the convergence speed and accuracy of the model by adaptively changing the learning rate based on the progress of the training. There are different strategies for dynamic learning rate, such as reducing the learning rate by a factor when the loss function stops improving, or increasing the learning rate when the model’s performance has plateaued. By adjusting the learning rate dynamically, the model can better navigate the loss landscape and converge to the global minimum more efficiently.
  • Hyperparameter tuning. It consists of strategies to choose the set of hyperparameters of the neural network to obtain good training. The most classic example is a hyperparameter grid, where all possible combinations of the hyperparameters to be determined are evaluated and the combination with the best performance of the loss function is chosen.
@@ III.2 Genetic algorithms
To apply genetic algorithms to a specific problem, one must select the objective function to optimize, delineate the search space, and specify the genetic parameters such as crossover, mutation, and elitism. Probability values for crossover and mutation operators are assigned, and a selection operator determines which individuals advance to the subsequent generation. Elitism, represented by a positive integer value, dictates the number of individuals guaranteed passage to the next generation. Overall, genetic algorithms initialize a population and iteratively modify individuals through operators and the objective function, progressively approaching the optimal solution of the target function.
@@ IV Machine learning strategies
In this section, we outline our proposed method, which integrates machine learning techniques to implement neural networks and genetic algorithms within a nested sampling framework. Initially, we revisit the strategies enumerated in Section III.1, elucidating their application:
  • Data scaling. All samples within the parameter space are scaled to the interval $(0,1]$. We avoid zero to mitigate potential issues with rectified linear unit (ReLU) activation functions, which could lead to misleading information. Standardizing all parameter ranges to the same scale ensures consistency across parameters.
  • Early stopping. Early stopping not only allows to stop when the training does not improve; it also guarantees to keep the weights of the neural network that generated a lower value of the loss function, regardless of whether this configuration was from the last epoch or a previous one. We use early stopping, because we want a low number of epochs around $100$, but we are interested in keeping the best weights.
  • Dynamic learning rate. There are popular strategies for dynamic learning rates. However, our dynamic learning rate is only adjusted during the nested sampling run and not during the training of a specific neural network. For each new training of the neural network, the learning rate decreases by half.
  • Hyperparameter tunning. We have implemented the option of using genetic algorithms to find the architecture of the first trained neural network. For this purpose, we use the library nnogada gomez2023neural .
We implemented our method inside of the code SimpleMC simplemc ; aubourg2015  [^1] The modified version of SimpleMC that includes our neuralike method is available at https://github.com/igomezv/simplemc_tests/tree/neuralike, which uses the library dynesty speagle2020dynesty for nested sampling algorithms.
@@
Neural networks are widely acclaimed for their formidable capabilities in handling extensive datasets. However, several studies have shown their effectiveness in modeling small datasets as well; even demonstrating that neural models can accommodate a total number of weights exceeding the number of sample data points ingrassia2005neural . Recent research has focused on novel approaches to using neural networks with smaller datasets ng2015deep ; pasini2015artificial ; gomez2023neuralepjc . In nested sampling, as discussed in the previous section, there is a set of live points that maintain a constant number of elements. At a certain point in its execution, a new sample is extracted within a prior iso-likelihood or mass surface. Our goal is for the neural network to predict the likelihood of points within this prior volume. To do this, we train the neural network with only the current set of live points. These points, which typically are around the hundreds or thousands, are sufficient to effectively train a neural network and have several advantages:
@@
  • By excluding points outside the current prior volume, we can potentially avoid inaccurate predictions in regions where points would be rejected based on the original likelihood. The points that the neural network learns efficiently are those within the prior volume, which have a higher probability of acceptance according to the original likelihood.
@@ V Toy models
We have used some toy models as log-likelihood functions: Gaussian, egg-box, and Himmelblau. In Table 1,1, you can see the results of the Bayesian evidence calculation with and without our method for the three toy models, while in Figure 1, you can see the samples of the three functions, which at first glance are very similar. Based on these results, we can notice that for all these models, the speed of sampling using neural networks is slower than in the case of nested sampling alone; this is because the analytical functions are being evaluated directly without sampling from an unknown posterior distribution; nevertheless, these examples are very useful to verify the accuracy in calculating Bayesian evidence and sampling from the distribution. We can observe that both the log-Bayesian evidence and the graphs of the nested sampling process without and with neural networks are consistent; however, as Table 1 shows, for more complex functions, we need a lower value of dlogz_start, which means that we need to start learning the neural network at a later stage of nested sampling. Therefore, a lower dlogz_start parameter is needed to be more accurate but slower, and it is precisely this parameter that regulates the speed-accuracy trade-off.
@@ VI Cosmological parameter estimation
  • Planck-15 information. We also consider a compressed version of Planck-15 information, where the Cosmic Microwave Background (CMB) is treated as a BAO experiment located at redshift $z=1090$, measuring the angular scale of the sound horizon. For more details, please see the Reference aubourg2015 .
We executed three cases of parameter estimation to verify the performance of our method. We start with one thousand live points and a model with five free parameters; then, we increase the live points and free parameters, to test our method with higher dimensionality and with higher computational power demand (larger number of live points). The results are compared with a nested sampling run with the same data sets and the same configuration (live points, stopping criterion, etc.) but without ANN; this comparison aims to test the accuracy and speedup achieved by our neuralike method. For this comparison, we report the parameter estimation and Bayesian evidence obtained with and without our method and, in addition, we calculate the Wasserstein distances ramdas2017wasserstein between the posterior nested sampling samples without and with neuralike for each free parameter considering their respective sampling weights.
@@ VI.2 Case 2
Secondly, we consider the same model, free parameters, and datasets as in Case 1. The difference in this second case is to analyze the behavior of our method with a larger number of live points. It has three new considerations: a) the training set for the neural network would be better because has a larger size, b) the number of operations in parallel for nested sampling is also larger, and c) we test the hypotheses based on a larger number of live points can obtain a better accuracy for the neural network earlier within the nested sampling process (i. e. in a higher value for dlogz_start). Therefore, we increase the number of live points to 4000 and dlogz_start = 20; the outputs are included in Figure 3 showing an excellent concordance for the Bayesian evidence values with our method, and speed-up around the $28.4\%$. Table 2 contains the results of the Bayesian evidence, and it can be noticed that the uncertainty of this case is in better agreement with nested sampling than the two scenarios of Case 1. In addition, we can analyze Table 3 and conclude that, effectively, its performance has a similar quality to Case 1 with dlogz_start = 5; however because it uses a higher dlogz_start value, the percentage of saved time is notorious.

### v2

@@ Abstract
In this paper, we present a novel approach to accelerate the Bayesian inference process, focusing specifically on the nested sampling algorithms. Bayesian inference plays a crucial role in cosmological parameter estimation, providing a robust framework for extracting theoretical insights from observational data. However, its computational demands can be substantial, primarily due to the need for numerous likelihood function evaluations. Our method utilizes the power of deep learning, employing feedforward neural networks to approximate the likelihood function dynamically during the Bayesian inference process. Unlike traditional approaches, our method trains neural networks on-the-fly using the current set of live points as training data, without the need for pre-training. This flexibility enables adaptation to various theoretical models and datasets. We perform the hyperparameter optimization using genetic algorithms to suggest initial neural network architectures for learning each likelihood function. Once sufficient accuracy is achieved, the neural network replaces the original likelihood function. The implementation integrates with nested sampling algorithms and has been thoroughly evaluated using both simple cosmological dark energy models and diverse observational datasets. Additionally, we explore the potential of genetic algorithms for generating initial live points within nested sampling inference, opening up new avenues for enhancing the efficiency and effectiveness of Bayesian inference methods.
@@ I Introduction
Bayesian inference is a powerful tool in several scientific fields where it is essential to constrain mathematical models using experimental data. It allows parameter estimation and model comparison. In particular, it is the data analysis technique per excellence in observational cosmology, as it provides a robust method to obtain valuable statistical information from a theoretical model given a set of observational data. However, a significant disadvantage of Bayesian inference lies in its high computational cost; it requires a considerable number of likelihood function evaluations to generate sufficient samples from the posterior distribution. For example, a small Bayesian inference task could involve thousands of samples and require thousands, or even millions, of likelihood evaluations.
@@
The use of artificial neural networks (ANNs) to approximate the likelihood function can greatly improve the efficiency of Bayesian inference auld2007fast ; graff2012 ; graff2014 ; hortua2020accelerating ; hortua2020parameter ; spurio2022cosmopower ; nygaard2023connect . However, it is necessary to have a careful consideration of the trade-off between accuracy and speed, along with quality monitoring of the resulting posterior samples. In addition, neural networks present several drawbacks that must be taken into account to effectively aid in the performance of Bayesian inference:
@@
Previous works using neural networks in cosmological parameter estimation save an amazing amount of computational time training neural networks before the Bayesian inference process spurio2022cosmopower ; hortua2020accelerating ; chantada2023nn ; however, the pre-training time in these cases is expensive and the trained neural networks are only useful for a specific configuration of backgrounds, models, and data sets. For this reason, our work is inspired by BAMBI graff2012 ; graff2014 , and pyBambi pybambi , where their neural networks are trained in real-time to learn the likelihood function, which is subsequently replaced within a nested sampling process. The strength of this approach lies in its ability to train the neural network in real-time and accelerate the Bayesian inference process without restricting a cosmological or theoretical model and specific datasets. In our method, we explore features beyond those of our predecessors, such as parallelism, PyTorch implementation paszke2019pytorch , and hyperparameter tuning. In addition, we exclusively used live points for training to reduce the dispersion of the training dataset and to obtain results with higher accuracy. A criterion was also chosen to initiate our method that serves as a regulator of the trade-off between accuracy and speed. We also implemented an on-the-fly performance evaluation to accept or reject the neural network predictions. In addition, we have conducted a preliminary investigation on the use of genetic algorithms to generate the initial sample of live points on the nested sampling process.
@@ II.1 Bayesian inference
and if the error $\epsilon\sim N(0,C)$ has a normal distribution centered in zero and a covariance matrix $C$, then we have the following:
$P(D|\theta)=\frac{1}{(2\pi)^{N/2}|C|^{1/2}}e^{-0.5(D-f(x;\theta))^{T}C^{-1}(D-f(x;\theta))}\;\;.$ (4)
@@ II.2 Nested sampling
NS starts with a specific number $n_{\rm live}$ of random points, termed live points, distributed within the prior volume defined by the constrained prior. These samples are ordered based on their likelihood values. During each iteration the worst point $\mathcal{L}_{\rm worst}$, with the lowest likelihood value, is removed. A new sample is then generated within a contour bounded by $\mathcal{L}_{\rm worst}$ and with a likelihood, $\mathcal{L}(\theta)>\mathcal{L}_{\rm worst}$. Equation 7 can be simplified as a Riemann sum:
@@
where $\omega_{i}$ is the difference between the prior volume of two consecutive points: $\omega_{i}=X_{i-1}-X_{i}$. Throughout the process, NS retains the population of $n_{\rm live}$ live points and ultimately consolidates the final set of live points within a region of high probability. Depending on the sampling approach employed from the constrained prior, various nested sampling algorithms exist. For instance, MultiNest feroz2009multinest utilizes rejection sampling within ellipsoids, whereas Polychord handley2015 generates points using slice sampling.
@@ III Machine Learning background
Machine learning is the field of Artificial Intelligence concerning to the mathematical modeling of datasets. Its methods identify inherent properties of datasets by minimizing a target function until it reaches a satisfactory value. Over the past few years, Artificial Neural Networks (ANNs) have emerged as the most successful type of machine learning models, giving rise to the field of deep learning. On the other hand, genetic algorithms are a special class of evolutionary algorithms, called metaheuristics, facilitating function optimization without derivatives.
@@ III.1 Artificial neural networks
The intrinsic parameters of a neural network, known as hyperparameters, are set before training, and include parameters such as the number of layers and neurons, epochs, and activation functions. Parameters of gradient descent and backpropagation algorithms rumelhart1986learning , like batch size and learning rate, may also be hyperparameters. While some hyperparameters are predetermined, others are adjusted through tuning strategies.
ANNs are valued for their capacity to model large and complex datasets. The Universal Approximation Theorem asserts that an ANN with a single hidden layer and non-linear activation functions can model any nonlinear function hornik1990universal , enhancing its utility for datasets with complex relationships. Even though an exhaustive review of ANNs is beyond the scope of this paper, great references exist in the literature nielsen2015neural ; goodfellow2016deep . For a basic introduction to their algorithms in the cosmological context, we recommend reading de2022observational .
@@ III.2 Genetic algorithms
To apply genetic algorithms to a specific problem, one must select the objective function to optimize, delineate the search space, and specify the genetic parameters such as crossover, mutation, and elitism. Probability values for crossover and mutation operators are assigned, and a selection operator determines which individuals advance to the subsequent generation. Elitism, represented by a positive integer value, dictates the number of individuals guaranteed passage to the next generation. Overall, genetic algorithms initialize a population and iteratively modify individuals through the operators and the objective function, progressively approaching the optimal solution of the target function.
@@ IV Machine learning strategies
In this section, we outline our proposed method, which integrates machine learning techniques to implement neural networks and genetic algorithms within a nested sampling framework. Below we describe some deep learning techniques utilized in our training:, elucidating their application:
  • Data scaling. Since all samples within the parameter space are already scaled between 0 and 1 during nested sampling, no additional scaling is required for training the neural networks.
  • Early stopping. It is a regularization technique that monitors the performance of a model on a validation set during training and stops the training process when the performance on the validation set starts to degrade, indicating overfitting. It helps to prevent overfitting and choose the best weight configuration along the epochs of the training. By stopping the training process early, the generalization performance of the model can be improved, particularly when the training data is limited or noisy. We implement early stopping with a patience of 100 epochs to guarantee a minimum number of training epochs, given the smaller size of the dataset. However, our primary focus is on preserving the best-performing weights at the end of the training process.
  • Dynamic learning rate. There are popular strategies for dynamic learning rates. However, our dynamic learning rate is only adjusted during the nested sampling run and not during the training of a specific neural network. For each new training of the neural network, the learning rate decreases by half. However, during each individual ANN training session, the learning rate remains constant within the adaptive gradient descent algorithm called Adam kingma2014adam .
  • Hyperparameter tunning. We have implemented the option of using genetic algorithms to find the architecture of the first trained neural network. For this purpose, we use the library nnogada gomez2023neural . For simplicity in this work, we use genetic algorithms over 3 generations with a population size of 5 to explore combinations of batch size (4 or 8), number of layers (2 or 3), learning rate (0.0005 or 0.001), and number of neurons per layer (50 or 100). In a nested sampling execution, where we can train the neural networks multiple times, we use these small configurations. This approach yields better results compared to not tuning hyperparameters and is more effective than using a hyperparameter grid gomez2023neural .
We implemented our method inside of the code SimpleMC simplemc ; aubourg2015  [^1] The modified version of SimpleMC that includes our neuralike method is available at https://github.com/igomezv/simplemc_tests, which uses the library dynesty speagle2020dynesty for nested sampling algorithms. In all our neural network training, we use the mean squared error (MSE) as the loss function. If early stopping, with a patience of 100 epochs, does not stop the training, we select the configuration of weights that achieved the lowest MSE value.
@@
Neural networks are widely acclaimed for their formidable capabilities in handling extensive datasets. However, several studies have shown their effectiveness in modeling small datasets as well; even demonstrating that neural models can accommodate a total number of weights exceeding the number of sample data points ingrassia2005neural . In addition, recent research has focused on novel approaches by using neural networks with smaller datasets ng2015deep ; pasini2015artificial ; gomez2023neuralepjc . While it is true that models with a large number of parameters can be prone to overfitting, this risk can be mitigated through the use of regularization techniques such as dropout and early stopping. In our approach, these techniques, combined with genetic algorithms for optimizing the network’s architecture and hyperparameters, ensure that our models generalize well even when the number of parameters exceeds the number of data points.
In nested sampling, as discussed in the previous section, there is a set of live points that maintain a constant number of elements. At a certain point in its execution, a new sample is extracted within a prior iso-likelihood or mass surface. Our goal is for the neural network to predict the likelihood of points within this prior volume. To do this, we train the neural network with only the current set of live points. These points, which typically are around hundreds or thousands, are sufficient to effectively train a neural network and have several advantages:
@@
  • By excluding points outside the current prior volume, we can potentially avoid inaccurate predictions in regions where points would be rejected based on the original likelihood. The points that the neural network learns efficiently are those within the prior volume, becasue they have a higher probability of acceptance according to the original likelihood.
@@
The likelihood function in cosmological parameter estimation can be quite complex, often involving various types of observational data and intricate numerical operations, such as integrals, derivatives, or approximation methods for solving differential equations. To address this complexity, the idea is to replace the analytical likelihood function with a trained neural network. This substitution reduces the problem to a simple matrix multiplication, where the optimal weights, obtained during ANN training, are stored in a binary file. Consequently, the evaluation of the likelihood becomes significantly faster. This acceleration is particularly advantageous in a Bayesian inference process, where the likelihood function may need to be evaluated thousands or even millions of times, making the reduction in computational time highly beneficial.
@@
It is important to note that the nested sampling process, including the selection of priors, typically uniform or Gaussian distributions, remains consistent with standard practices. Once the criteria for initiating ANN training are met, the live points are used to train the ANN. If the ANN’s performance metrics meet the required threshold, the analytical likelihood function is replaced by the ANN to save computational time. While this substitution does not alter the fundamental nested sampling process, it can significantly enhance efficiency by reducing computational overhead.
@@ V Toy models
We have used some toy models as log-likelihood functions: Gaussian, egg-box, and Himmelblau. In Table 1, you can see the results of the Bayesian evidence calculation with and without our method for the three toy models, while in Figure 1, you can see the samples of the three functions, which at first glance are very similar. Based on these results, we can notice that for all these models, the speed of sampling using neural networks is slower than in the case of nested sampling alone; this is because the analytical functions are being evaluated directly without sampling from an unknown posterior distribution; nevertheless, these examples are very useful to verify the accuracy in calculating Bayesian evidence and sampling from the distribution. We can observe that both the log-Bayesian evidence and the graphs of the nested sampling process without and with neural networks are consistent; however, as Table 1 shows, for more complex functions, we need a lower value of dlogz_start, which means that we need to start learning the neural network at a later stage of nested sampling. Therefore, a lower dlogz_start parameter is needed to be more accurate but slower, and it is precisely this parameter that regulates the speed-accuracy trade-off.
@@ VI Cosmological parameter estimation
  • Planck-15 information. We also consider a compressed version of Planck-15 information, where the Cosmic Microwave Background (CMB) is treated as a BAO experiment located at redshift $z=1090$, measuring the angular scale of the sound horizon. For more details, see the Reference aubourg2015 .
We executed three cases of parameter estimation to verify the performance of our method. We start with one thousand live points and a model with five free parameters; then, we increase the live points and free parameters, to test our method with higher dimensionality and with higher computational power demand (larger number of live points). The results are compared with a nested sampling run with the same data sets and the same configuration (live points, stopping criterion, etc.) but without ANN; this comparison aims to test the accuracy and speedup achieved by our neuralike method. For this comparison, we report the parameter estimation and Bayesian evidence obtained with and without our method and, in addition, we calculate the Wasserstein distances ramdas2017wasserstein between the samples of the posterior nested sampling without and with neuralike for each free parameter considering their respective sampling weights.
@@ VI.2 Case 2
Secondly, we consider the same model, free parameters, and datasets as in Case 1. The difference in this second case is to analyze the behavior of our method with a larger number of live points. It has three new considerations: a) the training set for the neural network would be better because has a larger size, b) the number of operations in parallel for nested sampling is also larger, and c) we test whether the hypotheses based on a larger number of live points can obtain a better accuracy for the neural network earlier within the nested sampling process (i.e. in a higher value for dlogz_start). Therefore, we increase the number of live points to 4000 and dlogz_start = 20; the outputs are included in Figure 3 showing an excellent concordance for the Bayesian evidence values with our method, and speed-up around the $28.4\%$. Table 2 contains the results of the Bayesian evidence, and it can be noticed that the uncertainty of this case is in better agreement with nested sampling than the two scenarios of Case 1. In addition, we can analyze Table 3 and conclude that, effectively, its performance has a similar quality to Case 1 with dlogz_start = 5; however because it uses a higher dlogz_start value, the percentage of saved time is notorious.